OpenStudy (anonymous):

A 0.4 ml dose of a drug is injected into a patient steadily for half a second. At the end of this time, the quantity, Q, of the drug in the body starts to decay exponentially at a continuous rate of 0.5% per second. Using formulas, express Q as a continuous function of time, t in seconds. Q=___________ for 0 ≤ t ≤ 0.5 and Q=___________ for 0.5 < t Thanks :)

OpenStudy (anonymous):

for the first one, Q=___________ for 0 ≤ t ≤ 0.5, i'm not sure... would it be either (4/5) t ? or 0.4t ? :/ and i'm kinda lost on the second one. Q=___________ for 0.5 < t i'm thinking that it starts off with 0.4e^ of something.. :/ but I'm not even sure if thats right haha :/

OpenStudy (agent0smith):

The initial amount is 0.4, the 0.5% per second means that after each second, 99.5% remains. Convert that to a decimal an... Q = 0.4(0.995)^t

OpenStudy (agent0smith):

for this Q=___________ for 0 ≤ t ≤ 0.5 it's injected steadily, ie linearly. 0.4 mL for 0.5 seconds... means 0.4/0.5 mL/second, or 4/5*t

OpenStudy (anonymous):

ohh okay :) so this equation, Q = 0.4(0.995)^t, is for 0.5 < t ?

OpenStudy (anonymous):

ohh that's what i mixed up haha so the first one is (0.4/0.5)t ?

OpenStudy (agent0smith):

0.4/0.5 = 4/5

OpenStudy (anonymous):

okay awesome!! thanks :)

OpenStudy (anonymous):

okay awesome!! thanks :)